A screw was detached from a hot air balloon that is rising at 3.5 m/s at an altitude of 8.3 m. (a) What is the initial velocity of the screw? (b) How long will it take for the screw to reach the ground?
1
Expert's answer
2020-10-07T10:13:26-0400

(a) the initial velocity of the screw is directed upright

"v_i=3.5\:\rm m/s"

(b) the equation of motion of the screw

"y=8.3+3.5t-\frac{9.8t^2}{2}"

When the screw reaches the ground "y=0," so

"8.3+3.5t-4.9t^2=0"

"t=1.7\:\rm s"